Deepa did a mathematics test with a differt design.
Each correct answer scored 8marks and each incorrect answer reduced the score by 4 marks. The test contained 30 questions,and after completing it deepa had a score of zero. How many questions did she answered correctly?

Answer by ankor@dixie-net.com(22740)

Deepa did a mathematics test with a different design.
Each correct answer scored 8 marks and each incorrect answer reduced the score by 4 marks.
The test contained 30 questions, after completing it Deepa had a score of zero.
:
Poor Deepa, some days it just doesn't pay to get out of bed.
:
Let x = no. of right answers
then
(30-x) = no. of wrong answers
:
8x - 4(30-x) = 0
8x - 120 + 4x = 0
8x + 4x = 120
12x = 120
x = 120/12
x = 10 right answers
:
:
See of that checks out: (She got 20 question wrong)
8(10) - 4(20) =
80 - 80 = 0